# Orchestrated Robust Controller for the Precision Control of Heavy-duty Hydraulic Manipulators

### v1

@@ II-B Lemmas and Assumptions
Because only the friction between the piston and cylinder in hydraulic actuators is critical, friction moments in all the rotating joints are equal to zero.
@@
[…]
$Z(\chi)=\hat{W}^{T}\Psi(\chi)+\hat{\varepsilon}$
[…]
where $\Xi_{N}=\{\hat{W}|\lVert\hat{W}\rVert\leq\kappa\}$ is a valid set of vectors with $\kappa$ being a design value, $\Xi_{T}$ is an allowable set of the state vectors, and $\hat{Z}(\chi|\hat{W})=\hat{W}^{T}\,\Psi(\chi)$.
Lemma 2
[36, 38]. For any inertial parameter vector $\phi_{A}$, there is a one-to-one linear map $f:\mathbb{R}^{10}\rightarrow S(4)$ such that,
$f(\phi_{A})=\mathcal{L}_{A}=\begin{bmatrix}0.5tr(\bar{I}).\textbf{1}-\bar{I}&h\\ h^{T}&m\end{bmatrix}$
$f^{-1}(\phi_{A})=\phi_{A}(m,h,tr(\Sigma).\textbf{1}-\Sigma)$
where $\Sigma=0.5tr(\bar{I})-\bar{I}$, and $m$,$h$, and $\bar{I}$ are the mass, first mass moment, and rotational inertia matrix, respectively, with $\phi_{A}=[m,h,vecI]$ where $vecI=[\bar{I}_{xx},\bar{I}_{yy},\bar{I}_{zz},\bar{I}_{xy},\bar{I}_{yz},\bar{I}_{xz}]$. Moreover, $tr(.)$ is the Trace operator of a matrix.
Lemma 3
[38]. For $\mathcal{L}_{A}$ defined in Lemma (2), Bregman divergence with the log-det function can be defined as,
$\mathcal{D}_{F}(\mathcal{L}_{A}\rVert\hat{\mathcal{L}}_{A})=log\frac{|\hat{\mathcal{L}}_{A}|}{|\mathcal{L}_{A}|}+tr(\hat{\mathcal{L}}_{A}^{-1}\mathcal{L}_{A})-4.$
@@
$\dot{\mathcal{D}}_{F}(\mathcal{L}_{A}\rVert\hat{\mathcal{L}}_{A})=tr([\hat{\mathcal{L}}_{A}^{-1}\dot{\hat{\mathcal{L}}}_{A}\,\hat{\mathcal{L}}_{A}^{-1}]\,\tilde{\mathcal{L}}_{A})$
where ~$\tilde{\mathcal{L}}_{A}=\hat{\mathcal{L}}_{A}-\mathcal{L}_{A}$. The Bregman divergence denotes the distance between the actual value $\mathcal{L}_{A}$ and its estimation ^$\hat{\mathcal{L}}_{A}$ over the manifold $\mathcal{M}\simeq\{\mathcal{L}_{A}\in\mathcal{S}(4):\mathcal{L}_{A}\succ 0\}=\mathcal{P}(4)$, with $\mathcal{S}(4)$ being space of $4\times 4$ real-symmetric matrices and $\mathcal{P}(4)$ being space of $4\times 4$ real-symmetric, positive-definite matrices.
@@ III Modeling the Rigid body subsystem
In this section, equations of motion for the 6-DoF HHM are derived. Since VDC is an NE-based approach, it is required to perform forward velocity computation and force backpropagation to achieve dynamic equations. As demonstrated in Fig. 1b, the entire complex system is divided into three objects, which are analyzed separately. Such a decomposition may help ease the understanding of the modeling and control design, as the HHM in this study is a complex system consisting of a rack and pinion joint, serial-parallel joints, and hydraulic rotary joints. By employing VDC, each subsystem, representing a rigid body or actuator, is treated individually.
@@ III-A1 Kinematics of Object 1
Consider ${{}^{G}V}$ to be the linear/angular velocity vector of the ground. The pillar velocity ${{}^{P_{1}}V}$ can be computed according to Fig. 2:
@@
with ${{}^{G}{U_{P_{1}}}}$ being in the sense of (2). In (8), the $\dot{\zeta}_{1}$ is the angular velocity of the pillar that can be computed by numerically differentiating the encoder angle data. This angular motion is generated by the piston’s linear motion through the rack and pinion mechanism. For a more accurate analysis, the dynamics of the piston must be considered as well. According to Fig. 2, we have:
${{}^{P_{p2}}V}={{}^{P_{p1}}{U_{P_{p2}}^{T}}}\,{{}^{P_{p1}}V}+x_{f}\dot{x}_{p}$ (9)
with $\dot{x}_{p}$ being the linear velocity of the piston. The following relation between the angular velocity of the pillar and the linear velocity of the piston holds:
@@ III-A2 Kinematics of Object 2
As demonstrated in Fig. 1b, object 2 is composed of two closed-chain kinematics, each of which has important geometric relations that represent constraints in the chain. In the following, the mentioned geometric relations are first expressed, and then linear/angular velocity vectors are derived. In the notation, $j=1,2$ represents the first and second closed chains, respectively. Both chains have a passive revolute joint driven by a linear hydraulic actuator. Hence, we have:
@@
$q_{j2}=-\arccos\left(\dfrac{L_{j}^{2}-(x_{j}+x_{j0})^{2}-L_{j1}^{2}}{-2\,(x_{j}+x_{j0})\,L_{j1}}\right),$ (14)
where $L_{j}$ and $L_{j1}$ are lengths between the frames, and $x_{j0}$ is the effective length of the hydraulic linear actuator at zero piston stroke. Differentiating (12)–(14) and appropriately transforming them result in:
@@
with $\dot{q}_{1}=\dot{\zeta}_{2}$ and $\dot{q}_{2}=\dot{\zeta}_{3}$. The equations in (11)-(17) reveal the relations between passive joint angle motion and corresponding piston motion. Through these relations, one can control the revolute joint angle by controlling the piston position.
Fig. 3: VDC frames of object 2
Exploiting VCP, closed chains in object 2 are decomposed into subsystems (shown in Fig. 3) through open chains, such that the linear/angular velocities in the open kinematic chain $1j$ and $2j$ can be written as:
@@
Considering Fig. 3, linear/angular velocities at the driving VCP of the closed kinematic chains are:
@@
Note that, from Fig. 3, we have:
@@
Finally, linear/angular velocities measured and expressed in frame $\left\{B_{c3}\right\}$ of object 2 can be calculated as:
@@
Therefore, all the linear/angular velocity vectors of object 2 are computed.
@@ III-A3 Kinematics of Object 3
Object 3 consists of three rotary hydraulic actuators (RHAs), as demonstrated in Fig. 4. For the sake of the compact computation, $i=1,2,3$ attributes to the first, second, and third rotary actuators, respectively. Accordingly, considering Fig. 4, we have:
@@
with ${{}^{E_{1}}V}={{}^{B_{c3}}V}$, and $\varrho_{1,3}={x}_{f}$ and $\varrho_{2}={y}_{f}$. Moreover, ${}^{E}V={{}^{E_{4}}{U}_{E}^{T}}\,{{}^{E_{4}}}V$ is the linear/angular velocity of the end-effector. It must be mentioned that the angular velocity of each joint of object 3 has the following relation with piston displacement inside the hydraulic rotary actuator:
@@ III-B Dynamics of the Manipulator
In this section, the total force/moment vectors exerted on each rigid body in each object are derived. The net forces in (4) can be computed for each rigid body by replacing $\left\{A\right\}$ with the corresponding frame. In order to consider the unknown disturbance and unknown model uncertainty, we can rewrite (4) as:
$M_{A}\dfrac{\rm d}{\mathrm{d}t}\left({{}^{A}V}\right)+C_{A}\left({{}^{A}{\omega}}\right){{}^{A}V}+G_{A}+\,{{}^{A}}\Delta_{R}(t)={{}^{A}F^{*}}-\,{{}^{A}}D(t)$ (31)
where ${{}^{A}}\Delta_{R}(t)$ and ${{}^{A}}D(t)$ satisfy Assumption 1 and being unmodeled dynamics and disturbances, respectively.
@@ III-B1 Dynamics of Object 3
The net force/moment vector of object 3 in Fig. 4 can be expressed as:
@@
for $i=3,2,1$. The ${{}^{E_{4}}F}={{}^{{E_{4}}}{U}_{E}}\,{{}^{E}F}$ with ${{}^{E}F}$ is the force/moment sensed from the environment. Consequently, the linear piston force in RHAs can be computed as:
@@ III-B2 Dynamics of Object 2
where ${{}^{{B_{c{3}}}}F}={{}^{{E_{1}}}F}$. Considering (36)-(40) and provoking theorems 3 and 4 of [6], the force/moment vector at the driven VCP along with linear actuator force of the parallel-serial joints in object 2, can be derived as:
@@ III-B3 Dynamics of Object 1
Having ${{}^{B_{c1}}F}$ from object 2 and considering Fig. 3, one can obtain the force/moment vector of the pillar as:
@@
${{}^{P_{p2}}F^{*}}$ is the net force/moment vector of the piston body.
@@ IV Control Design of the rigid body subsystem
In this section, for the modeled 6-DoF HHM, the procedure of local controller design is expressed. The VDC approach enables us to design local controllers that accomplish the local goals, and then by employing VPF, the main objective of the system is achieved by orchestrating all the local controllers.
@@ IV-A Desired Trajectory
To determine the precise motion of HHMs in Cartesian space, it is critical to have a robust and precise controller at the actuator level. The kinematic relation between the Cartesian space and joint space for the 6-DoF manipulator can be written as:
@@
with $\dot{\Pi}_{d}=[\dot{X}_{d},\,\dot{Y}_{d},\,\dot{Z}_{d},\,\dot{\varpi}_{xd},\,\dot{\varpi}_{yd},\,\dot{\varpi}_{zd}]^{T}$ being desired pose velocity of the end-effector, $\dot{\Theta}_{d}=[\dot{\zeta}_{1d},\dot{\zeta}_{2d},\dot{\zeta}_{3d},\dot{\xi}_{1d},\dot{\xi}_{2d},\dot{\xi}_{3d}]^{T}$ being desired angular velocity, and $J\in\mathbb{R}^{6\times 6}$ is the Jacobian matrix. Therefore, the joint space velocity can be computed as:
@@
It can be seen from (49) that the desired trajectory of working space can be projected onto a joint space trajectory. It means that precise joint control will end up with precise end-effector pose control. This study aims to design a precise controller to achieve a considerably low tracking error.
@@ IV-B Required Velocity Vectors
By solving the inverse kinematic problem in (49), the desired angular velocity and desired joint angle can be computed. Then, the required angular velocity in the sense of (6) can be derived. As expressed in (10) and (30), the linear motion is converted to angular motion through a constant parameter. However, the expression in (15) is complex and depends on the kinematics of the closed chain. Consequently, we define the required velocity of objects 1 and 3 based on the angular motion, and for object 2 based on piston motion. It is straightforward to compute the desired piston velocity and position for object 2 using (49), (11), (12), and (15) by replacing the actual signals with desired ones. Therefore, the required joint velocities can be defined as:
@@
with $j=1,2$, $i=1,2,3$, and $\lambda$, $\bar{\lambda}_{xj}$, and $\sigma_{i}$ are constant positive gains. By replacing (50)-(52) in the kinematic part of each object and reusing (10), (15)-(17), and (30), the required linear/angular velocities can be achieved as:
@@
The required joint velocities in (50)-(52) show the required trajectory that each joint must follow in order to achieve the local control objectives. On the other side, the required linear/angular velocity vectors in (53)-(64) display the velocity that the rigid body should acquire to track the desired motion. In the following, the local controller will be designed in a way that provides the required effort to establish the desired motion.
@@ IV-C Required Force Vectors
The required force/moment vector must be robust in the presence of model uncertainty and external disturbances. Therefore, considering the presented model in (31), the required net force/moment vector can be designed in the sense of (7) by taking advantage of (5) as:
@@
However, the model uncertainty ${{}^{A}}\Delta_{R}$ is unknown in real-world applications. Additionally, the magnitude and frequency of the external disturbances are unknown for many applications, especially for industrial manipulators. In this study, adaptive control and DRBFNN approaches are utilized to address the mentioned issues. According to Lemma 1, we can define:
@@
with ${{}^{A}}\hat{\varepsilon}$ being the estimation of ${{}^{A}}\varepsilon$. In (67), the term ${{}^{A}}\hat{\varepsilon}$ estimates both the bias of the RBFNNs ${{}^{A}}\varepsilon^{*}$ and the external disturbance $D(t)$. Then, the required force/moment vectors can be computed by evoking force/moment vectors and using (67) as:
@@
In the following section, taking into account the complex and nonlinear model of the hydraulic actuator, the low-level controller is designed to ensure the required forces in (74)-(76) will be generated at the actuator level. Since the procedure of modeling for all the hydraulic actuators in this study is the same, we derived the required voltage for a given linear piston force $f_{c}$, similar to (35), (42), and (47), and required piston force $f_{cr}$ as in (74)-(76).
@@ V-A Inverse Deadzone-Backlash
Input constraints must be considered in control design to ensure control signals remain within safe and feasible bounds. Ignoring input constraints can lead to issues, such as system instability, equipment damage, and safety hazards, especially in real-world applications. The compound input constraint, that is deadzone and backlash, is examined in this section. Suppose that $u_{d}^{*}$ is a control signal that may accomplish control objectives in the absence of input constraints, while $v^{*}$ is the control signal with deadzone-backlash compensation. The deadzone function can be defined as,
@@
with $m_{d}>0$, $b_{r}>0$, and $b_{l}<0$ being deadzone parameters. Backlash function can be defined as,
@@
with $k_{b}>0$, $B_{r}>0$, and $B_{l}<0$ being backlash parameters. By substituting (77) in (78), one can obtain,
@@
After designing the control voltage in the absence of compound input constraint (79), the control law can be enhanced to cancel the damaging effect of such a constraint, such that $u_{d}^{*}=DB(DBI(u_{d}^{*}))$ with $DBI(.)$ being the inverse deadzone-backlash function. On the other hand, in real-world applications, the parameters in (79) are unknown, and, it is desirable to estimate them. Consequently, we can write the adaptive inverse deadzone-backlash control term as,
@@
with $\eta_{0}$ being the bound mismatch error, $u^{*}(t)$ being the signal after deadzone-backlash, $u_{d}^{*}(t)$ being the desired signal in the absence of deadzone-backlash, and $\tilde{\theta}$ being the estimation error of the constraint parameters. To guarantee robustness and better performance, the parameter and model uncertainties of the compound input constraint can be tackled by designing an adaptation law for $\hat{\theta}$ and RBFNNs for estimation of $\eta_{0}$, which will be addressed later. Fig. 5 displays the closed-loop scheme of the original VDC and the proposed method. As it is shown in Fig. 5b the deadzone-backlash compensator is incorporated into the controller to assure robustness and better performance in comparison to the original VDC, shown in Fig. 5b.
Fig. 5: a) The closed-loop scheme of the original VDC in the presence of compound input constraint and unknown model and actuator uncertainties, b) proposed method with deadzone-backlash and uncertainty compensator
@@
The deadzone-backlash compensation error in (85) is general for any system that is subjected to such a compound constraint. Therefore, it can be incorporated into any control law at the force/torque and voltage levels to cancel the input constraint.

### v2

@@ II Mathematical preliminaries
Friction moments in all the rotating joints are neglected.
Definition 6
[37]. RBFNNs can be utilized to estimate an unknown continuous function $Z(\chi):\mathbb{R}^{\textit{m}}\to\mathbb{R}$ with the approximation of,
$Z(\chi)=\hat{W}^{T}\Psi(\chi)+\hat{\varepsilon}$
The optimal weight vector $W^{*}$ can be expressed by,
where $\Xi_{N}=\{\hat{W}|\lVert\hat{W}\rVert\leq\kappa\}$ is a valid set of vectors with $\kappa$ being a design value, $\Xi_{T}$ is an allowable set of the state vectors, and $\hat{Z}(\chi|\hat{W})=\hat{W}^{T}\,\Psi(\chi)$. Additionally, $\hat{\varepsilon}$, being the estimation error, is a function of input data.
Remark 1
While deep neural networks (DNNs) can model complex patterns by stacking multiple layers, this often requires significant computational resources and careful tuning. Since inn real-time systems where low latency and fast processing are critical, RBFNNs are often a better choice than DNNs.
Definition 7
[36, 38]. For any inertial parameter vector $\phi_{A}$, there is a one-to-one linear map $\mathcal{T}:\mathbb{R}^{10}\rightarrow S(4)$ such that,
$\mathcal{T}(\phi_{A})=\mathcal{L}_{A}=\begin{bmatrix}0.5tr(\bar{I}).\textbf{1}-\bar{I}&h\\ h^{T}&m\end{bmatrix}$
$\mathcal{T}^{-1}(\mathcal{L}_{A})=\phi_{A}(m,h,tr(\Sigma).\textbf{1}-\Sigma)$
where $\Sigma=0.5tr(\bar{I})-\bar{I}$ and $tr(.)$ is the Trace operator of a matrix.
@@
[38]. For $\mathcal{L}_{A}$ defined in Definition (7), Bregman divergence with the log-det function can be defined as,
$\mathcal{D}_{F}(\mathcal{L}_{A}\rVert\hat{\mathcal{L}}_{A})=log\frac{|\hat{\mathcal{L}}_{A}|}{|\mathcal{L}_{A}|}+tr(\hat{\mathcal{L}}_{A}^{-1}\mathcal{L}_{A})-4,$
@@
$\dot{\mathcal{D}}_{F}(\mathcal{L}_{A}\rVert\hat{\mathcal{L}}_{A})=tr([\hat{\mathcal{L}}_{A}^{-1}\dot{\hat{\mathcal{L}}}_{A}\,\hat{\mathcal{L}}_{A}^{-1}]\,\tilde{\mathcal{L}}_{A}).$
The Bregman divergence denotes the distance between the actual value $\mathcal{L}_{A}$ and its estimation ^$\hat{\mathcal{L}}_{A}$ over the manifold $\mathcal{M}\simeq\{\mathcal{L}_{A}\in\mathcal{S}(4):\mathcal{L}_{A}\succ 0\}=\mathcal{P}(4)$, with $\mathcal{S}(4)$ being space of $4\times 4$ real-symmetric matrices and $\mathcal{P}(4)$ being space of $4\times 4$ real-symmetric, positive-definite matrices.
@@ III Modeling the Rigid body subsystem
In this section, equations of motion for the generic 6-DoF HHM are derived. As demonstrated in Fig. 1b, the entire complex system is divided into three objects, which are analyzed separately. Such a decomposition may help ease the understanding of the modeling and control design, as the HHM in this study is a complex system consisting of a rack and pinion joint, serial-parallel joints, and hydraulic rotary joints. By employing VDC, each subsystem, representing a rigid body or actuator, is treated individually.
@@ III-A1 Kinematics of Object 1
Consider ${{}^{G}V}$ to be the spatial velocity vector of the ground. The pillar velocity ${{}^{P_{1}}V}$ can be computed according to Fig. 2:
@@
with ${{}^{G}{U_{P_{1}}}}$ being in the sense of (2), and $\dot{\zeta}_{1}$ is generated by the piston’s linear motion through the rack and pinion mechanism. According to Fig. 2, the piston spatial velocity can be obtained as:
${{}^{P_{p2}}V}={{}^{P_{p1}}{U_{P_{p2}}^{T}}}\,{{}^{P_{p1}}V}+x_{f}\dot{x}_{p},$ (9)
indicating the following relation between the angular velocity of the pillar and the linear velocity of the piston:
@@ III-A2 Kinematics of Object 2
As demonstrated in Fig. 1b, object 2 is composed of two closed-chain kinematics, each of which has important geometric relations that represent constraints in the chain. In the following, the mentioned geometric relations are first expressed, and then spatial velocity vectors are derived. In the notation, $j=1,2$ represents the first and second closed chains, respectively. Both chains have a passive revolute joint driven by a linear hydraulic actuator, as shown in Fig. 3b. Hence, we have:
@@
$q_{j2}=-\arccos\left(\dfrac{L_{j}^{2}-(x_{j}+x_{j0})^{2}-L_{j1}^{2}}{-2\,(x_{j}+x_{j0})\,L_{j1}}\right).$ (14)
Differentiating (12)–(14) and appropriately transforming them result in:
@@
with $\dot{q}_{1}=\dot{\zeta}_{2}$ and $\dot{q}_{2}=\dot{\zeta}_{3}$. The equations in (11)-(17) reveal the relations between angular motion of passive joints and corresponding piston motion.
Fig. 3: a) VDC frames of object 2, b) Closed chain mechanism with hydraulic actuator
Exploiting VCP, closed chains in object 2 are decomposed into subsystems (shown in Fig. 3a) through open chains, such that the spatial velocities in the open kinematic chain $1j$ and $2j$ can be written as:
@@
Considering Fig. 3a, spatial velocities at the driving VCP of the closed kinematic chains are:
@@
Note that, from Fig. 3a, we have:
@@
Finally, spatial velocities measured and expressed in frame $\left\{B_{c3}\right\}$ of object 2 can be calculated as:
@@
Therefore, all the spatial velocity vectors of object 2 are computed.
@@ III-A3 Kinematics of Object 3
According to Fig. 4, we have:
@@
with $i=1,2,3$, ${{}^{E_{1}}V}={{}^{B_{c3}}V}$, and $\varrho_{1,3}={x}_{\tau}$ and $\varrho_{2}={y}_{\tau}$. Moreover, ${}^{E}V={{}^{E_{4}}{U}_{E}^{T}}\,{{}^{E_{4}}}V$ is the spatial velocity of the end-effector. Additionally, the following relation between angular motion and piston displacement holds:
@@ III-B Dynamics of the Manipulator
In this section, the total spatial force vectors exerted on each rigid body in each object are derived. The net forces in (4) can be computed for each rigid body by replacing $\left\{A\right\}$ with the corresponding frame. In order to consider the unknown disturbance and unknown model uncertainty, we can rewrite (4) as:
$M_{A}\dfrac{\rm d}{\mathrm{d}t}\left({{}^{A}V}\right)+C_{A}\left({{}^{A}{\omega}}\right){{}^{A}V}+G_{A}+\,{{}^{A}}\Delta_{R}(t)={{}^{A}F^{*}}-\,{{}^{A}}D(t).$ (31)
@@ III-B1 Dynamics of Object 3
The net spatial force vector of object 3 in Fig. 4 can be expressed as:
@@
for $i=3,2,1$. The ${{}^{E_{4}}F}={{}^{{E_{4}}}{U}_{E}}\,{{}^{E}F}$ with ${{}^{E}F}$ is the spatial force sensed from the environment. Consequently, the linear piston force in RHAs can be computed as:
@@ III-B2 Dynamics of Object 2
where ${{}^{{B_{c{3}}}}F}={{}^{{E_{1}}}F}$. Considering (36)-(40) and provoking theorems 3 and 4 of [6], the spatial force vector at the driven VCP along with linear actuator force of the parallel-serial joints in object 2, can be derived as:
@@ III-B3 Dynamics of Object 1
Having ${{}^{B_{c1}}F}$ from object 2 and considering Fig. 3, one can obtain the spatial force vector of the pillar as:
@@
${{}^{P_{p2}}F^{*}}$ is the net spatial force vector of the piston body.
@@ IV Control Design of the rigid body subsystem
In this section, for the modeled 6-DoF HHM, the procedure of local controller design is expressed, where the main objective of the system can be achieved by orchestrating all the local controllers.
@@ IV-A Desired Trajectory
The kinematic relation between the Cartesian space and joint space for the 6-DoF manipulator can be written as:
@@
which results in:
@@
It can be seen from (47) that the desired trajectory of working space can be projected onto a joint space trajectory. It means that precise joint control will end up with precise end-effector pose control. This study aims to design a precise controller to achieve a considerably low tracking error in joint space.
@@ IV-B Required Velocity Vectors
By solving the inverse kinematic problem in (47), the desired angular velocity and desired joint angle can be computed. Then, the required angular velocity in the sense of (6) can be derived. As expressed in (10) and (30), the linear motion is converted to angular motion through a constant parameter.
However, the expression in (15) is complex and depends on the kinematics of the closed chain. Consequently, we define the required velocity of objects 1 and 3 based on the angular motion, and for object 2 based on piston motion. It is straightforward to compute the desired piston velocity and position for object 2 using (47), (11), (12), and (15) by replacing the actual signals with desired ones. Therefore, the required joint velocities can be defined as:
@@
with $j=1,2$, $i=1,2,3$. By replacing (48)-(50) in the kinematic part of each object and reusing (10), (15)-(17), and (30), the required spatial velocities can be achieved as:
@@
The required joint velocities in (48)-(50) show the required trajectory that each joint must follow in order to achieve the local control objectives. On the other side, the required spatial velocity vectors in (51)-(62) display the velocity that the rigid body should acquire to track the desired motion. In the following, the local controller will be designed in a way that provides the required effort to establish the desired motion.
@@ IV-C Required Force Vectors
Considering the presented model in (31), the required net spatial force vector can be designed in the sense of (7) by taking advantage of (5) as:
@@
However, the model uncertainty ${{}^{A}}\Delta_{R}$ is unknown in real-world applications. Additionally, the magnitude and frequency of the external disturbances are unknown for many applications, especially for industrial manipulators. In this study, adaptive control and DRBFNN approaches are utilized to address the mentioned issues. According to Definition 6, we can define:
@@
with ${{}^{A}}\hat{\varepsilon}$ being the estimation of ${{}^{A}}\varepsilon$. In (65), the term ${{}^{A}}\hat{\varepsilon}$ estimates both the bias of the RBFNNs ${{}^{A}}\varepsilon^{*}$ and the external disturbance $D(t)$. Then, the required spatial force vectors can be computed by evoking spatial force vectors and using (65) as:
@@
In the following section, by taking into account the complex and nonlinear model of the hydraulic actuator, the low-level controller is designed to ensure the required forces in (72)-(74) will be generated at the actuator level. Since the procedure of modeling for all the hydraulic actuators in this study is the same, we derived the required voltage for a given linear piston force $f_{c}$, similar to (35), (42), and (45), and required piston force $f_{cr}$ as in (72)-(74).
@@ V-A Inverse Deadzone-Backlash
Suppose that $u_{d}^{*}$ is a control signal that may accomplish control objectives in the absence of input nonlinearities, while $v^{*}$ is the control signal with deadzone-backlash compensation. The deadzone function can be defined as,
@@
Backlash function can be defined as,
@@
By substituting (75) in (76), one can obtain,
@@
After designing the control voltage in the absence of compound input nonlinearities (77), the control law can be enhanced to cancel the damaging effect of such a constraint, such that $u_{d}^{*}=DB(DBI(u_{d}^{*}))$ with $DBI(.)$ being the inverse deadzone-backlash function. On the other hand, in real-world applications, the parameters in (77) are unknown, and, it is desirable to estimate them. Consequently, we can write the adaptive inverse deadzone-backlash control term as,
@@
To guarantee robustness and better performance, the parameter and model uncertainties of the compound input nonlinearities can be tackled by designing an adaptation law for $\hat{\theta}$ and RBFNNs for estimation of $\eta_{0}$, which will be addressed later. Fig. 5 displays the closed-loop scheme of the original VDC and the proposed method. As it is shown in Fig. 5b the deadzone-backlash compensator is incorporated into the controller to assure robustness and better performance in comparison to the original VDC.
Fig. 5: a) The closed-loop scheme of the original VDC in the presence of compound input nonlinearities and unknown model and actuator uncertainties, b) proposed method with deadzone-backlash and uncertainty compensator
@@
The deadzone-backlash compensation error in (83) is general for any system that is subjected to such a compound constraint. Therefore, it can be incorporated into any control law at the force/torque and voltage levels.